Home‎ > ‎Crew 135‎ > ‎

202-CommanderReport

Commander Report
02/02/2014
Ondrej Doule

The rest of the Crew134 was transported to Hanksville in the morning
after additional briefing on
science lab, ATVs, spacesuits and water refill process. Legal documents
"The Mars Society: MDRS Crew
Members Waiver, Release and Indemnity" were dropped in the mailbox in
Hanksville in the morning.
Digital copy of the documents is provided to all crew members as
required by the document.

Crew 135 spent all day by preapration of the research, simulation,
training with ATVs, systems and
equipment check Current systems state:

Spacesuits:
- Old backpack PLSS (6x): 1,2,3 are OK and charging; 4 does not charge
and the charging cable is
broken; 5,6 are OK but are not charging due to broken charging cables.
- New backpack PLSS (4x): 9,10,12,(no number)are not charging
- Old helmets (5x): 4x are OK; 1 is smashed
- New helmets (4x): 3x are OK with little cracks, 1 is broken
- Radios (7x): 5x is working OK and are charging, 2 are broken (FCCIDs:
K7GMJBGJ, K7GEM1000)
- Radio headset (4x): 1x is usable, 3x are not compatible (not usable)
- Suits: 4x L are OK, 4x M (1 with hole and 1 dirty) OK, 4x S OK but 3x
are very dirty
- Gloves: 5x skiing gloves OK (2x L, 1xM, 2x S (very dirty))
- Legging: only 2 pairs were found after 1h search randomly spread
around the airlock area (dirty).

ATVs:
- 1,2,3 OK, start well, full tanks,4 OK takes time to warm up and start
- two cannisters of petrol are approx halfull, one was placed on the
trailer one was next barrels of
which content we do not know.

Habs SUV:
- Odometer: 131643
- "Service engine soon" control is active
- Windshield has small spider crack (1 inch) in the center (probably
from a rock)
- Windshield, glasses, head/back lights were washed
- Tank was filled (was 3/4 empty)
- Interior requires very serious cleaning (will be done later)
- Gearshifting from P to R/D is a bit "loose"
- Parked in designed area behind power generator

According to HALs quick guides:
- Are Web cameras applicable to us_? we do not see any cameras plugged
in the hab or around.
- White water system - all OK and working according to guides.
- Gray water system - OK but N/A
- Black water system - Page in guides failed to open - we have no
instructions except from Shannon
(first telecon)
- EVA suit maintenance - OK
- EVA planning - OK currently using individual system / HALp system - we
are working on it

We are starting simulation tomorrow with experiments
- EXP01 (Critacality systems survey) inclusive 2h EVA (if approved)
- EXP02 (Greenhouse research)
- EXP05 (EVA helmet ergonomics) merged with EXP06
- EXP06 (MDRS growth - surroundings survey) inclusive afternoon 3h EVA
(if approved)
- EXP07 (Nutrition)